shift summary, day 21 June 99

Marj Corcoran (corcoran@physics.rice.edu)
Mon, 21 Jun 1999 16:05:39 -0500 (CDT)

Shift Summary for DAY shift, 21-June-99

On shift: J. Hamm, M. Corcoran, C. Wolfe

Hours of beam delivered at all intensitites: 5.0
Hours of beam delivered at nominal intensity: 5.0
Hours of E832 data (special runs not included): 5.0
Hours of special runs (laser scans, muon runs...): 0.5 hrs

Accounting of lost time: Beam off due to magnet problems

==================================================================

Runs ended this shift:
Run 13675: E832 data. Ended when beam went away to do laser scan
Run 13676: Laser scan, aborted when beam returned
Run 13677: E832 data, ended during down time to drt_fresh_start in
an (unsuccessful) attempt to fix DAQ screen which is not
updating spill
Run 13678: E832 data, ended when beam went off at 15:15.

Tapes brought to FCC: None

===================================================================

Description of shift:

Came in to find no beam due to power supply problem. After one hour and
no estimate for beam, we start a laser scan. Abort 2/3 of the way through
as beam returns.

8:50 AM Beam returns

Note on run statistics summary sheets: The proper way to check these
numbers is to log onto atev1 as sdaq, then "setup sdaq test".
Then "sdaq_online" should work.

NM2EU still not being read during spill by sdaq, so it comes up as an
error on the sdaq screen. The magnet is ramping and can be seen on the
EPICURE screen to be OK.

10:30 Beam goes away due to magnet problems.

11:00 Rick Ford calls to say recalibration of SEM will occur tomorrow
(June 22) at 7AM. We also expect results of rad survey sometime tomorrow.

11:45 beam returns, resume data-taking

12:00 Rick T calls to alert us to a problem with HA channel 28. It
has a very low ADC spectrum on the last muon run. We look at the signal
going into the ADC interface and notice that the negative-going part of the
signal looks normal, but it is followed by a large positive overshoot
which is not present on any of the other HA signals. We conclude that
the ADC is subtracting the positive overshoot, resulting in a near-zero
net signal. The signal going into the ADC interface also shows the
positive overshoot, so the problem is downstairs. We would require an access
to investigate further.

12:45 Note that mean SODs are a bit high in the current run, max of
0.645. The sigmas are OK. Muon monitoring reports possible timing shift
which may account for slight shift in sods.

15:10 MCR calls, beam will go away in 5 min for an hour.

15:30 MDC and CW make an access to look at HA28. It appears that the problem is
in the shaper/splitter, since the signal coming into it looks fine and
exactly like Ch 27. So need to find a spare channel for HA28.

===================================================================

Followup notes on previous shift:

KTeV DAQ window is still not updating spill number and run number, and
run summaries are not getting written. We did a drt_fresh_start during
the beam down time, but it did not solve the problem. Arun suggests
logging out, opening a new window, and then doing a drt_fresh_start if
the opportunity presents itself.

===================================================================
Items to follow up on next shift:

KTeV DAQ window still not updating. Try opening a new window before
doing drt_fresh_start.